Differentiate f(x) = (5x³+7x )/(2x²-4x+5 ) using quotient rule.
1
Expert's answer
2020-09-13T18:27:00-0400

"f'(x)=\left(\frac{5x^3+7x}{2x^2-4x+5}\right)'=\\
{}=\frac{(5x^3+7x)'(2x^2-4x+5)-(5x^3+7x)(2x^2-4x+5)'}{(2x^2-4x+5)^2}=\\
{}=\frac{(15x^2+7)(2x^2-4x+5)-(5x^3+7x)(4x-4)}{(2x^2-4x+5)^2}=\\
{}=\frac{10x^4-40x^3+61x^2+35}{(2x^2-4x+5)^2}."